OpenStudy (anonymous):

what is 2.89 Square Root. Please i need Solution. THANKS!

OpenStudy (saifoo.khan):

can u draw it?

OpenStudy (anonymous):

actually the answer is 1.7 but i need solution.

myininaya (myininaya):

\[\sqrt{289}=17\] \[\sqrt{289 \cdot \frac{1}{100}}=\frac{\sqrt{289}}{\sqrt{100}}=\frac{17}{10}\]

OpenStudy (anonymous):

ok Thank you!